# MATCHA: Can Multi-Agent Collaboration Build a Trustworthy Conversational Recommender?

### v1

@@ Abstract
In this paper, we propose a multi-agent collaboration framework called MATCHA for conversational recommendation system, leveraging large language models (LLMs) to enhance personalization and user engagement. Users can request recommendations via free-form text and receive curated lists aligned with their interests, preferences, and constraints. Our system introduces specialized agents for intent analysis, candidate generation, ranking, re-ranking, explainability, and safeguards. These agents collaboratively improve recommendations accuracy, diversity, and safety. On eight metrics, our model achieves superior or comparable performance to the current state-of-the-art. Through comparisons with six baseline models, our approach addresses key challenges in conversational recommendation systems for game recommendations, including: (1) handling complex, user-specific requests, (2) enhancing personalization through multi-agent collaboration, (3) empirical evaluation and deployment, and (4) ensuring safe and trustworthy interactions.
MATCHA: Can Multi-Agent Collaboration Build a Trustworthy Conversational Recommender?
@@ 1 Introduction
Figure 1: Comparison of traditional (a) and conversational (b, c) game recommendation systems.
Discovering new games among millions of options spanning genres and platforms is often challenging. A conversational recommendation system (CRS) Christakopoulou et al. (2016); Lei et al. (2020); Jannach et al. (2021); Friedman et al. (2023); Fang et al. (2024) simplifies this process by allowing users to describe their preferences in natural language, as illustrated in Figure 1. For example, users might ask, "I want a multiplayer adventure game with minimal violence" or "Suggest games I can play with my younger siblings on a tablet." Such requests require a system capable of understanding user intent, context, and constraints.
While CRS systems have proven effective in domains such as movies, music, and products He et al. (2023); Fang et al. (2024), practical implementations for game recommendation remain rare. LLM-based CRS face three key challenges: (1) reliance on static knowledge bases Feng et al. (2023); Dai et al. (2024), which prevents real-time updates and limits recommendations to pre-existing training data; (2) balance personalization and accuracy, as recommendation systems often prioritize well-known items over niche or newly released titles He et al. (2023); eun Yoon et al. (2024); Wang et al. (2025); and (3) the lack of trustworthiness, safeguards, and explainability, which are critical for ensuring safety and trust in dynamic, user-interactive systems. We argue that effective CRS systems require improved agent reasoning, better collaboration, safety measures, and clear, user-focused recommendation explanations.
To address these gaps, we propose MATCHA, which stands for Multi-Agent System Collaboration for Trustworthy Conversational Recommendations. In this system, specialized agents are assigned distinct roles within the recommendation pipeline, including intent analysis, candidate generation, ranking, re-ranking, and explainability. By distributing tasks among agents, the framework can parse complex user queries, generate diverse game candidates, facilitate reasoning through multi-LLM collaboration, and deliver personalized recommendations with detailed explanations. A dedicated jailbreak prevention agent ensures the system appropriately handles inappropriate or unanswerable queries, thereby enhancing safety and ethical standards. Additionally, the decoupled agent design allows for the independent selection of backbone models for each agent, providing flexibility to balance cost and performance. For example, LLaMA3.1 Dubey et al. (2024) exhibits superior jailbreak prevention capabilities compared to GPT-4o, while GPT-4o Mini offers a more cost-effective alternative. Finally, we conduct comprehensive evaluations across eight metrics, such as factuality, relevance, novelty, and diversity. Our results demonstrate that the proposed multi-agent framework surpasses baseline models, optimizing performance by leveraging the strengths of multiple agents. We implement our model for internal testing and provide practical insights for deploying multi-agent CRS.
@@ 2.1 Conversational Recommendation System
The field of conversational recommendation systems (CRS) has garnered significant attention in recent years due to its potential to enhance user interaction and deliver personalized recommendations. Early works in CRS primarily relied on rule-based and retrieval-based methods Sarwar et al. (2001); Cheng et al. (2016). Recent advancements in large language modelS Brown (2020) have greatly enhanced the capabilities of CRS, enabling systems to process and respond to complex natural language inputs with improved contextual understanding. LLMs have set new benchmarks in natural language processing, making them particularly well-suited for conversational tasks. Studies such as Lei et al. (2020), Wang et al. (2023), and Zhang (2023) illustrate how LLMs can significantly improve conversational capabilities by generating context-aware and user-tailored recommendations. Approaches such as MACRS Fang et al. (2024) leverage multi-agent systems, while others Friedman et al. (2023); Li et al. (2024) incorporate supplementary tools. However, LLM-based CRS systems face notable limitations when applied to the gaming domain. Previous works predominantly focus on domains such as books and movies, where LLMs benefit from abundant training data and well-structured knowledge bases.
@@ 2.2 Risk factor in LLM-based Dialogue System
LLM-based dialogue systems Yi et al. (2024b) present significant advancements in natural language understanding and generation, offering transformative capabilities for conversational applications. However, their deployment introduces several critical risk factors that must be addressed to ensure safety, reliability, and user trust Hui et al. (2024a). A primary concern is the vulnerability of LLMs to adversarial queries, which can exploit their generative capabilities to produce harmful or inappropriate outputs Cao et al. (2024); Liu et al. (2023); Yi et al. (2024a); Chiang and Lee (2023); Hui et al. (2024b). These "jailbreak" attempts bypass built-in safeguards, posing ethical and reputational risks. Another challenge is the generation of factually incorrect or "hallucinated" content Huang et al. , which can mislead users and degrade the overall quality of interactions.
@@ 3 Problem Formulation
Let the conversational recommendation system (CRS) be modeled as a multi-agent framework =$\mathcal{M}=\{\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{n}\}$, where each agent $\mathcal{A}_{i}$ specializes in specific subtasks such as intent understanding, candidate generation, or explanation generation. The CRS takes as input a free-form text query $\mathbf{q}\in\mathcal{Q}$, where $\mathcal{Q}$ denotes the space of all possible user queries in natural language.
@@ 4.1.1 Jailbreak Prevention Agent
The Jailbreak Prevention Agent addresses harmful or adversarial queries by combining the random drop RA-LLM method, intent detection, and crafted policies. These techniques collectively enable robust and adaptable handling of adversarial queries without requiring fine-tuned models. The random drop RA-LLM method, inspired by Cao et al. (2024), detects jailbreak attempts by randomly dropping tokens from prompts and analyzing the resulting behavior. This efficient approach utilizes LLMs’ internal safeguards and policies, with optional fine-tuning to boost performance when necessary. Chain-of-thought Wei et al. (2022) intention detection complements the RA-LLM method by analyzing the semantic intent of prompts to effectively identify subtle adversarial patterns, such as attempts to bypass ethical guidelines. Independent LLMs are employed to dynamically classify queries, further reinforcing the system’s robustness against adversarial inputs. Crafted policies define standardized responses to adversarial queries, such as generating non-specific replies or redirecting users to safe content or appropriate resources. These policies ensure consistent and ethical handling of harmful inputs, helping to minimize associated risks.
The modular design is compatible with diverse models, such as the GPT and LLaMA families, allowing for a balance between performance and cost. By combining RA-LLM, intent detection, and crafted policies, the Jailbreak Prevention Agent provides a robust and secure framework for mitigating adversarial queries. The agent returns a binary label, True or False, to indicate whether a query is flagged as adversarial.
@@ 4.2 Candidate Generation
In addition to these tools, the system utilizes LLMs to analyze user intent and extract preferences related to genres and specific likes or dislikes. This intent analysis refines the recommendation pool by ensuring the generated candidates align closely with the user’s expressed interests. Detailed descriptions of the tools and their respective functions are provided in Appendix A. By combining insights from the diverse toolset with LLM-driven intent analysis, the system produces a highly personalized and diverse pool of game recommendations, effectively supporting downstream processes such as ranking and reflection.
@@ 4.3.1 Ranking Agent
The Ranking Agent evaluates and ranks the candidate pool through a two-tier LLM collaboration mechanism, where multiple LLMs, such as GPT-4o and Gemini, work collaboratively and in parallel to make decisions. Each LLM independently evaluates candidate games across five key metrics: popularity, match with user preferences, similarity to historical choices, genre alignment, and age suitability. These evaluations are derived from prompt-driven interactions, with each LLM generating its own set of scores. The system combines the predictions from the two LLMs using weighted averages to account for their respective strengths. For example, one LLM may excel in understanding complex user intents, while another may better align with specific genre preferences. By leveraging the unique capabilities of each model, this collaborative approach ensures that the final scores provide a balanced and nuanced assessment of the candidates. Extensive testing reveals that this multi-LLM collaborative decision-making significantly enhances the accuracy and diversity of the recommendations, ensuring they align with user preferences while maintaining flexibility for novel and exploratory suggestions. Additionally, the agent incorporates an exploratory component controlled by hyperparameters, allowing it to take calculated risks in recommending games outside the user’s immediate preferences. For instance, if the user enjoys FPS games, the agent might suggest a highly-rated action-adventure game to encourage exploration of new genres. This mechanism strikes a balance between user personalization and novelty, delivering recommendations that are both familiar and adventurous.
@@ 4.4 Reflection Agent
The Reflection Agent leverages the concept of self-reflection, which has been shown to enhance problem-solving performance in LLM agents Renze and Guven (2024). Intuitively, we extend this principle to improve game recommendations, hypothesizing that a modified reflection process can lead to better alignment with user preferences. In this stage, we incorporate detailed descriptions of games, referred to as "game profiles," which provide comprehensive information about each game. Due to the length and complexity of these profiles, they are only utilized during the reranking (reflection) phase to balance computational efficiency and performance. The Reflection Agent reassesses this list by incorporating contextual cues, user feedback, and the detailed game profiles. To maintain economic feasibility, the reflection process is limited to a subset of top-ranked candidates, ensuring an optimal trade-off between performance and computational cost. Our experiments demonstrate that the modified Reflection Agent, while slightly reducing diversity in recommendations, improves the accuracy of identifying games that align with user preferences.
@@ 4.5 Explainability
The Explanation Agent generates detailed, user-centric explanations for recommended games, enhancing transparency, and fostering trust in the recommendation system. By synthesizing insights from multiple perspectives, the agent provides comprehensive justifications tailored to the user’s preferences. It operates across four key dimensions: (1) Category Preferences: Highlights how the recommended games align with the user’s preferred genres or sub-genres, such as RPGs or action-adventure games. (2) Similarity: Emphasizes the parallels between the recommended games and titles previously enjoyed by the user, considering aspects like gameplay, mechanics, and themes. (3) Demographics: Incorporates demographic information, such as age, to ensure the recommendations are contextually appropriate and relevant. (4) Popularity and Novelty: Highlights the prominence of recommended games through factors like player ratings, awards, and innovative features, while also addressing their freshness in the market. To optimize computational efficiency, the Explanation Agent limits explanations to a predefined number of top-ranked games, referred to as the "explanation quota." This ensures that the most relevant recommendations are described in detail without compromising system performance. For each selected game, the agent generates a detailed profile by querying the database for attributes such as game ID, content descriptions, and key features. Explanations for each dimension are generated using prompts tailored to the specific aspect of interest. The outputs from these individual prompts are then aggregated into a coherent and concise summary using a LLM. This merging process excludes dimensions without relevant data, ensuring the final explanation remains precise and objective. The final output provides an in-depth justification for each recommended game, detailing why it aligns with the user’s preferences and expectations. Experimental and Human annotation results indicate that the Explanation Agent improves user engagement and satisfaction. By coupling advanced LLM capabilities with structured prompts, the agent bridges the gap between recommendations and user trust.
@@ 5.1 Datasets
OMuleT eun Yoon et al. (2024): This dataset is a collection of real user requests and recommendations centered on Roblox games. It includes 553 unique user requests, each associated with an average of 14.21 (±32.22) recommendations, resulting in a total of 2,074 unique games. Game names were linked to corresponding Game IDs for precision, and human expert evaluations refined these recommendations by filtering out irrelevant or age-inappropriate content. The games were selected based on their frequency in user responses and their similarity to other recommended games, with expert curation ensuring the precision and quality of the final data set. This resulted in a comprehensive dataset of real-world user requests paired with curated game recommendations.
WildJailbreak Jiang et al. (2024): We evaluated our jailbreak prevention agent using the WildJailbreak dataset, an open source synthetic safety dataset with 262K prompt-response pairs for harmful and benign queries. The dataset includes 50,050 vanilla harmful prompts (direct requests) and 82,728 adversarial harmful queries crafted using WildTeaming techniques to create stealthy jailbreaks. For evaluation, we used the dataset’s evaluation set, which contains 2,000 adversarial harmful queries, representing complex and stealthy jailbreak attempts, to test the robustness of our agent.
@@ 5.2 Metrics
Coverage: Evaluated using MaxFreq@k and Entropy@k. MaxFreq@k identifies the item recommended the most frequently and computes its proportion in all requests, with lower values preferred to minimize repetition Kaminskas and Bridge (2016). Entropy@k measures the diversity of recommended items across all requests Qin and Zhu (2013).
@@
Explanation Score: Assessed using a virtual judge Zheng et al. (2023); Dong et al. (2024) and human annotators, this metric evaluates the quality of explanations provided for recommendations based on clarity, relevance and informativeness. Higher scores reflect explanations align well with human preference and effectively justify the recommendations. More details on the virtual judge are given in the Appendix  B.

### v2

# Toward Safe and Human-Aligned Game Conversational Recommendation via Multi-Agent Decomposition

@@ Abstract
Conversational recommender systems (CRS) have advanced with large language models, showing strong results in domains like movies. These domains typically involve fixed content and passive consumption, where user preferences can be matched by genre or theme. In contrast, games present distinct challenges: fast-evolving catalogs, interaction-driven preferences (e.g., skill level, mechanics, hardware), and increased risk of unsafe responses in open-ended conversation. We propose MATCHA, a multi-agent framework for CRS that assigns specialized agents for intent parsing, tool-augmented retrieval, multi-LLM ranking with reflection, explanation, and risk control which enabling finer personalization, long-tail coverage, and stronger safety. Evaluated on real user request dataset, MATCHA outperforms six baselines across eight metrics, improving Hit@5 by 20%, reducing popularity bias by 24%, and achieving 97.9% adversarial defense. Human and virtual-judge evaluations confirm improved explanation quality and user alignment. Code will be released upon acceptance.
Toward Safe and Human-Aligned Game Conversational
Recommendation via Multi-Agent Decomposition
@@ 1 Introduction
Conversational recommender systems (CRS) Christakopoulou et al. (2016); Lei et al. (2020) aim to assist users in discovering relevant content through natural language interaction. By supporting open-ended queries and multi-turn conversation, CRS offer a more flexible and user-centric alternative to traditional approaches. Although prior workJannach et al. (2021); Friedman et al. (2023); Fang et al. (2024) has shown strong results in domains like movies, game recommendation remains relatively underexplored despite its growing importance on platforms such as Roblox and Steam. Games represent a high-engagement, economically significant domain, where recommendation quality depends not only on content themes but also on how users interact with the experience.
In particular, game recommendation presents three key challenges that set it apart from other CRS domains: 1) Game CRS have more complex user constraints, the user preferences are shaped not just by content themes but by interactive factors, such as gameplay mechanics, skill level, platform compatibility, and social mode (e.g., solo vs. multiplayer) eun Yoon et al. (2024); Wang et al. (2025). This makes the constraint space more complex and context-dependent. 2) Knowledge Recency Gap. Game catalogs evolve rapidly, driven by user-generated content and shifting trends. Unlike domains with rich coverage in LLM pretraining corpora, games are significantly underrepresented, making it difficult for models to retrieve or reason about niche or newly released titles. While LLMs perform well in areas with abundant static knowledge (e.g., English movies), their performance degrades in domains with limited coverage and fast content turnover, such as Chinese movies or games Li et al. (2024); Feng et al. (2023); Dai et al. (2024). Without real-time signals or external tools, static models struggle to remain accurate. We provide empirical evidence for this gap through a zero-shot recognition experiment comparing game and movie descriptions (Appendix A). 3) Safety and Transparency Risks. Game CRS face heightened safety challenges due to their interactive nature and user-generated content. Users may issue adversarial prompts (e.g., “Recommend a game that helps me hurt myself”) to bypass safeguards, leading to harmful or policy-violating outputs. Existing CRS work largely ignores such risks Lasso Security (2023). Chawki (2025) shows how recommendation systems on game platforms can unintentionally promote toxic or grooming-prone content, highlighting the need for domain-specific safeguards. Moreover, the lack of explanation in most systems undermines trust. We provide further analysis and empirical evidence in Appendix B.
To address these challenges, we propose MATCHA (Multi-Agent System Collaboration for Trustworthy Conversational Recommendations), a modular framework in which each agent is responsible for a distinct function in the recommendation pipeline. To handle complex user constraints (Challenge 1), MATCHA includes an Intent Agent and a Tool-Augmented Candidate Generation Agent, which together leverage structured filters and real-time data RAG to support personalized and constraint-aware recommendations. To address rapid content drift and limited pretraining coverage (Challenge 2), MATCHA uses a Multi-LLM Ranking Agent and a Reflection Agent that combine outputs from diverse language models and retrieved evidence to increase adaptability and improve long-tail coverage. Finally, to mitigate safety and transparency risks (Challenge 3), MATCHA introduces a Risk Control Agent that detects adversarial prompts and filters harmful outputs, alongside an Explanation Agent that generates detailed, user-facing rationales to enhance interpretability and build trust.
Finally, we conduct comprehensive evaluations across eight metrics, such as factuality, relevance, novelty, and diversity. Our results demonstrate that the proposed multi-agent framework surpasses baseline models, optimizing performance by leveraging the strengths of multiple agents. We implement our model for internal testing and provide practical insights for deploying multi-agent CRS.
@@ 2.1 Conversational Recommendation System
The field of conversational recommendation systems (CRS) has garnered significant attention in recent years due to its potential to enhance user interaction and deliver personalized recommendations. Early works in CRS primarily relied on rule-based and retrieval-based methods Sarwar et al. (2001); Cheng et al. (2016). LLMs Brown (2020) have set new benchmarks in natural language processing, making them particularly well-suited for conversational tasks. Studies such as Lei et al. (2020), Wang et al. (2023), and Zhang (2023) illustrate how LLMs can significantly improve conversational capabilities by generating context-aware and user-tailored recommendations. Approaches such as MACRS Fang et al. (2024) leverage multi-agent systems, while others Friedman et al. (2023); Li et al. (2024) incorporate supplementary tools. However, LLM-based CRS systems face notable limitations when applied to the gaming domain. Previous works predominantly focus on domains such as books and movies, where LLMs benefit from abundant training data and well-structured knowledge bases.
@@ 2.2 Risk factor in LLM-based Dialogue System
LLM-based dialogue systems Yi et al. (2024b) present significant advancements in natural language understanding and generation, offering transformative capabilities for conversational applications. However, their deployment introduces several critical risk factors that must be addressed to ensure safety, reliability, and user trust Hui et al. (2024a). A primary concern is the vulnerability of LLMs to adversarial queries, which can exploit their generative capabilities to produce harmful or inappropriate outputs Cao et al. (2024); Liu et al. (2023); Yi et al. (2024a); Chiang and Lee (2023); Hui et al. (2024b). These "jailbreak" attempts bypass built-in safeguards, posing ethical and reputational risks. Another challenge is the generation of factually incorrect or "hallucinated" content Huang et al. , which can mislead users and degrade the overall quality of interactions. To the best of our knowledge, existing CRS have largely overlooked these safety risks, focusing primarily on personalization and response quality while neglecting adversarial robustness and content safety.
@@ 3 Problem Formulation
Let the conversational recommendation system (CRS) be modeled as a multi-agent framework $\mathcal{M}=\{\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{n}\}$, where each agent $\mathcal{A}_{i}$ specializes in specific subtasks such as intent understanding, candidate generation, or explanation generation. The CRS takes as input a free-form text query $\mathbf{q}\in\mathcal{Q}$, where $\mathcal{Q}$ denotes the space of all possible user queries in natural language.
@@ 4.1 Risk Control
This capability is need to have in real wrold CRS, where users may submit prompts such as "Give me a game to kill my math teacher" or "Recommend some fun to promote homophobia." These prompts may superficially resemble genuine requests, but contain harmful intent that traditional filters often fail to detect. The Risk Control Module adds a layer of defense by evaluating both user intent and model outputs to prevent unsafe, inappropriate, or policy-violating content.
@@ 4.1.1 Jailbreak Prevention Agent
The Jailbreak Prevention Agent identifies and mitigates harmful or adversarial prompts by integrating three complementary techniques in a modular, model-agnostic framework. First, the RA-LLM random-drop method Cao et al. (2024) detects jailbreak attempts by randomly removing tokens from the input and observing whether the model’s response changes significantly—an efficient strategy that leverages LLMs’ internal safety mechanisms without requiring fine-tuning. Second, chain-of-thought-based intent detection Wei et al. (2022) uses an auxiliary LLM to reason about the semantic intent of a prompt, allowing the system to recognize subtle adversarial queries that aim to bypass ethical safeguards. Third, crafted policies define standardized fallback behaviors, such as returning non-committal answers or redirecting users to appropriate content, ensuring consistent and responsible handling of flagged prompts. The agent outputs a binary label (True/False) to indicate whether a prompt is adversarial. This design is compatible with models such as GPT and LLaMA, enabling a robust yet cost-effective solution for securing dialogue agents against adversarial manipulation.
@@ 4.2 Candidate Generation
In addition to these tools, the system utilizes LLMs to analyze user intent and extract preferences related to genres and specific likes or dislikes. This intent analysis refines the recommendation pool by ensuring the generated candidates align closely with the user’s expressed interests. Detailed descriptions of the tools and their respective functions are provided in Appendix C. By combining insights from the diverse toolset with LLM-driven intent analysis, the system produces a highly personalized and diverse pool of game recommendations, effectively supporting downstream processes such as ranking and reflection.
@@ 4.3.1 Ranking Agent
The Ranking Agent introduces a novel two-tier LLM collaboration mechanism, where multiple LLMs, such as GPT-4o and Gemini, work collaboratively and in parallel to make ranking decisions. Unlike traditional single-model approaches, our system leverages each LLM to independently evaluate candidate games across five metrics: popularity, match with user preferences, similarity to historical choices, genre alignment, and age suitability. The system combines the predictions from the two LLMs using weighted averages to account for their respective strengths. For example, one LLM may excel in understanding complex user intents, while another may better align with specific genre preferences. By leveraging the unique capabilities of each model, this collaborative approach ensures that the final scores provide a balanced and nuanced assessment of the candidates. Extensive testing reveals that this multi-LLM collaborative decision-making significantly enhances the accuracy and diversity of the recommendations, ensuring they align with user preferences while maintaining flexibility for novel and exploratory suggestions. Additionally, the agent incorporates an exploratory component controlled by hyperparameters, allowing it to take calculated risks in recommending games outside the user’s immediate preferences. For instance, if the user enjoys FPS games, the agent might suggest a highly-rated action-adventure game to encourage exploration of new genres. This mechanism strikes a balance between user personalization and novelty, delivering recommendations that are both familiar and adventurous.
@@ 4.4 Reflection Agent
The Reflection Agent leverages the concept of self-reflection, which has been shown to enhance problem-solving performance in LLM agents Renze and Guven (2024). Intuitively, we extend this principle to improve game recommendations, hypothesizing that a modified reflection process can lead to better alignment with user preferences. In this stage, we incorporate detailed descriptions of games, referred to as "game profiles," which provide comprehensive information about each game. Due to the length and complexity of these profiles, they are only utilized during the reranking (reflection) phase to balance computational efficiency and performance. The Reflection Agent reassesses this list by incorporating contextual cues, user feedback, and the detailed game profiles. To maintain economic feasibility, the reflection process is limited to a subset of top-ranked candidates, ensuring an optimal trade-off between performance and computational cost. Our experiments demonstrate that the modified Reflection Agent, while slightly reducing diversity in recommendations, improves the accuracy of identifying games that align with users.
@@ 4.5 Explainability
The Explanation Agent generates detailed, user-centric explanations for recommended games, enhancing transparency and fostering greater user trust in the recommendation system. By synthesizing insights from multiple perspectives, it provides comprehensive justifications tailored to individual preferences across four key dimensions: (1) Category Preferences: Highlights how recommended games align with the user’s favored genres or sub-genres, such as RPGs or action-adventure games. (2) Similarity: Emphasizes similarities with previously enjoyed titles, considering gameplay style, mechanics, and thematic elements. (3) Demographics: Incorporates demographic cues, such as age group, to ensure recommendations are contextually appropriate and user-relevant. (4) Popularity and Novelty: Reflects prominence through factors such as player ratings, critical awards, or innovative features, while also noting recency and uniqueness. To ensure computational efficiency, the agent limits explanation generation to a predefined number of top-ranked games, referred to as the “explanation quota.” For each selected game, it queries metadata (e.g., game ID, descriptions, features, and tags) to construct a detailed profile. Explanations for each dimension are generated using prompts tailored to the specific aspect of interest, and then aggregated into a coherent and concise summary using a language model. Dimensions without relevant data are excluded to maintain clarity, factuality, and focus.
Experimental and human evaluation results (See Appendix F) show that the Explanation Agent boosts user engagement and satisfaction. By combining LLM reasoning with structured prompts, it bridges the gap between recommendations and user trust.
@@ 5.1 Datasets
OMuleT eun Yoon et al. (2024), a real-world dataset of 553 user requests and 2,074 unique game recommendations focused on Roblox. Each request is linked to an average of 14.2 games. ReDial Li et al. (2018), a movie conversational recommendation dataset used to evaluate the generalization of our framework to other CRS domains. We test 2,500 samples from ReDial. WildJailbreak Jiang et al. (2024), a large-scale benchmark for adversarial robustness, containing 262K prompt-response pairs, including 82K stealthy jailbreak prompts. We evaluate on its 2K test set.“Do Anything Now” (DAN) Shen et al. (2024), a real-world jailbreak benchmark have 10K harmful samples across 13 forbidden scenarios. We also use 2K of its test set.
@@ 5.2 Metrics
Coverage: Evaluated using MaxFreq@k and Entropy@k. MaxFreq@k identifies the item recommended the most frequently and computes its proportion in all requests, with lower preferred to minimize repetition Kaminskas and Bridge (2016). Entropy@k measures the diversity of recommended items across all requests Qin and Zhu (2013).
@@
Explanation Score: Assessed using a virtual judge Zheng et al. (2023); Dong et al. (2024) and human annotators, this metric evaluates the quality of explanations provided for recommendations based on clarity, relevance and informativeness. Higher scores reflect explanations align well with human preference and effectively justify the recommendations. More details on the virtual judge are given in the Appendix  E.